# Revisiting the integral form of Gauss’ law for a generic case of electrodynamics

### v2

@@ III Conclusion
Our calculations are based on the Maxwell equations which don’t need any relativistic correction. Hence the final result (i.e. Eqn. (12)) on the time-dependent form of Gauss’ law is essentially correct from the relativistic point of view.
@@
The charges inside and outside the arbitrarily moving Gaussian surface are considered to be moving arbitrarily. While the uniformly moving charges result in no radiation, the accelerating charges inside and outside the Gaussian surface result in electromagnetic radiation. This radiation does not change the form of Gauss’ law as clear from Eqn. (12).
Finally, it should be mentioned that the differential equation for Gauss’ flux-integral (Eqn. (10)) obtained by us, is novel, simple-looking, and important for determining time-dependent Gauss’ flux-integral. The time-dependent form of Gauss’ flux-integral (Gauss’ law), i.e. Eqn. (12), obtained by us although distinguished with the form $q_{\text{in}}(t)=q_{\text{in}}^{(0)}+\triangle q_{\text{in}}^{(s)}(t,t_{0})$, is essentially not a new result because it exactly matches with (Eqn. (6)) with the form $q_{\text{in}}(t)=\int_{V(t)}\rho(\vec{r},t)\text{d}^{3}\vec{r}$. Our explicit derivation of Gauss’ flux-integral for the moving surface can, however, be directly useful in an undergraduate/postgraduate physics class for a better understanding of Gauss’ law for electrodynamics.
@@
Acknowledgement
S. Biswas acknowledges the partial financial support of the SERB (now ANRF), DST, Govt. of India under the EMEQ Scheme [No. EEQ/2023/000788]. We thank Mr. Bikram Keshari Behera (UoH, India) for helping us draw figure 1. Useful discussions with Dr. Abhiram Soori (UoH, India), Dr. N. Sri Ram Gopal (UoH, India), Prof. V. Subrahmanyam (UoH, India), and Prof. Koushik Ray (IACS, India) are gratefully acknowledged.

### v3

@@ III Conclusions
Our calculations are based on the Maxwell equations which don’t need any relativistic correction. Hence, the final result (i.e. Eqn. (12)) on the time-dependent form of Gauss’ law is essentially correct from the relativistic point of view.
Since the Gaussian surface is an imaginary surface, its motion does not affect the electric field. The flux integral of the same electric field, however, changes with time whenever there is a change in the total amount of stationary/moving charges inside the moving Gaussian surface.
@@
While the uniformly moving charges result in no radiation, the accelerating charges inside and outside the Gaussian surface result in electromagnetic radiation. This radiation does not, however, change the form of Gauss’ law as clear from Eqn. (12).
Finally, it should be mentioned that the evolution equation (i.e. Eqn. (10)) obtained by us is novel, elegant, and important for determining the time-dependence of Gauss’ flux integral in a generic non-static case. Here, the use of the convective derivative $\frac{D\rho}{Dt}=\frac{\partial\rho}{\partial t}+\vec{v}^{\prime}\cdot\gradient{\rho}$ Jackson1 of the scalar field $\rho(\vec{r},t)$ captures the effect of the expansion/contraction, not the deformation. Here, we haven’t separately considered any overall displacement of the Gaussian surface because the expansion/contraction and the deformation together may result in an overall displacement. However, why there is no effect of the deformation on the flux integrals (Eqns. (10) and (12)), can be explained within our method with Eqn. (II) as a starting point under the consideration of the convective derivative of the vector field $\vec{E}(\vec{r},t)$. Thus, our explicit derivation of the flux-integral for the moving surface can be directly useful in an undergraduate (as well as postgraduate) physics class for a better understanding of Gauss’ law for electrodynamics.
Acknowledgement
S. Biswas acknowledges the partial financial support of the SERB (now ANRF), DST, Govt. of India under the EMEQ Scheme [No. EEQ/2023/000788]. We thank Mr. Bikram Keshari Behera (UoH, India) for helping us draw figure 1. Useful discussions with Dr. Abhiram Soori (UoH, India), Dr. Sudipto Muhuri (UoH, India), Dr. N. Sri Ram Gopal (UoH, India), Prof. V. Subrahmanyam (UoH, India), Mr. Rhitabrata Bhattacharyya (KCDCC(P), WB, India), and Prof. Koushik Ray (IACS, India) are gratefully acknowledged. We thank the anonymous reviewer for his/her thorough review and highly appreciate his/her critical remarks which have significantly contributed to improving the quality of the article.